OpenStudy (anonymous):

if you subtract 5 from me and then subtract 5 again, you get 7

OpenStudy (akashdeepdeb):

Do it JUST like the previous one! :)

OpenStudy (akashdeepdeb):

@RamyaR Are you getting me? :)

OpenStudy (anonymous):

no

OpenStudy (akashdeepdeb):

JUST LIKE THIS PREVIOUS QUESTION YOU POSTED WHICH WAS SIMILAR! Okay let us say I am a number x [Like Chemical X] So if you subtract 8 from me that is x-8 And then subtract 2 from me again x -8 -2 you get 5 So writing this like Math --> x -8-2 = 5 x-10=5 x - 10 + 10 = 5 + 10 [ADD 10 TO BOTH SIDES] x = 15 Yayy! So I am 15! You subtract 8 from me and get 7! You subtract 2 more you get 5 [YIPPEE!] Understood? :)

OpenStudy (anonymous):

isn't it 15

OpenStudy (anonymous):

17

OpenStudy (akashdeepdeb):

Yes the answer for this one IS 17! :)